OpenStudy (anonymous):

Find the x- and y-intercepts for the following equation. −4x + 3y = 9

OpenStudy (anonymous):

The y-intercept is when x=0 -4(0) + 3y = 9 -> 0 + 3y = 9 -> y = 3 The x-intercept is when y=0 -4x + 3(0) = 9 -> -4x + 0 = 9 -> x = -9/4

OpenStudy (anonymous):

So, the y-intercept is (0, 3) and the x-intercept is (-9/4, 0)

OpenStudy (anonymous):

All good now?

OpenStudy (anonymous):

So for Find the x- and y-intercepts for the following equation. (Examples 1-3) y = 4x + 8 would x =-2,0 and y= 0,8?

OpenStudy (anonymous):

yes, you got it!